OpenStudy (llivingston2):

Of the following options, what could be a possible first step in solving the equation −7x − 5 = x + 3? adding 7x to both sides of the equation subtracting 5 from both sides of the equation adding x to both sides of the equation combining like terms, -7x+x=-6x

OpenStudy (welshfella):

which one of those would eliminate the term in x on one side of the equation?

OpenStudy (llivingston2):

adding x to both sides?

OpenStudy (welshfella):

No that would not do it left side term would change to -6x and right side would become 2x

OpenStudy (welshfella):

suppose we had -x on the left what would happen if we added x to it?

OpenStudy (llivingston2):

it would be a positive

OpenStudy (welshfella):

-x + x = what?

OpenStudy (llivingston2):

a positive x

OpenStudy (welshfella):

no -x + x = 0

OpenStudy (andres19):

to solving that equation you have to move the x to one side, when you move the x you have to change the sign and add to the other x. lets move the right x to the left side −7x − 5 = x + 3 -8x=8 when you move the x to the left its a -x when u add its a -8x and the same with the 5 after that you have to leave the x alone so you have to divide the -8 to the 8 like this 8/-8 and the answer is -1

OpenStudy (llivingston2):

idk :(

OpenStudy (welshfella):

in the question we have -7x on left side so in order to get rid of this what do we add?

OpenStudy (llivingston2):

we add it to both sides

OpenStudy (welshfella):

yes we add it to both sides but what do we add -7x + something must = zero

OpenStudy (llivingston2):

-7x and -7x cancel each other out

OpenStudy (welshfella):

right!!

OpenStudy (welshfella):

u got it

OpenStudy (llivingston2):

thank you soo much!!

OpenStudy (welshfella):

yw

OpenStudy (llivingston2):

is that your kitty?

OpenStudy (andres19):